# Maximizing Asynchronicity in Event-based Neural Networks

### v1

@@ 3.2.1 Multi-representation prediction
For MRP, we force the learned representation to predict multiple handcrafted (converted) event representations lagorce2016hots ; sironi2018hats ; innocenti2021temporal , such as event count (EC) and time surface (TS), to learn a more comprehensive representation. For events =$\mathcal{E}=\{e_{i}=(t_{i},x_{i},y_{i},p_{i})\},i=1,2,\dots$, we expect the asynchronous encoder $\mathcal{M}_{\bm{\theta}}$ parametrized by $\bm{\theta}$ to learn an event-by-event representation as $\mathcal{R}^{\bm{\theta}}_{i}=\mathcal{M}_{\bm{\theta}}(\{e_{j}\}_{j\leq i})$. Given $K$ handcrafted representations $\{\mathcal{R}^{k}\}_{k=1}^{K}$, we predict these representations from $\mathcal{R}^{\bm{\theta}}$ with specific neural network heads parameterized by $\bm{\Theta}=\{\bm{\theta}_{k}\}_{k=1}^{K}$. The objective is:
$\arg\max_{\bm{\theta},\bm{\Theta}}\mathbb{E}_{i}\prod_{k=1}^{K}p\left(\mathcal{R}^{k}\left(\{e_{j}\}_{j\leq i}\right)|\mathcal{M}_{\bm{\theta}}\left(\{e_{j}\}_{j\leq i}\right),\bm{\theta}_{k}\right).$ (6)
MRP is motivated by the insight that different handcrafted representations capture distinct raw event information aspects and provide visual data for various tasks Zubic_2023_ICCV . We anticipate that learning to predict multiple converted representations will yield a representation generalizable to diverse downstream tasks.
@@ 3.2.2 Next representation prediction
Inspired by auto-regressive pretraining of language models, where models learn by next-token prediction radford2018improving , NRP forces the encoder to predict handcrafted representations for a future time interval. For a future time window $\Delta T$, the objective of NRP is:
$\arg\max_{\bm{\theta},\bm{\Theta}^{\prime}}\mathbb{E}_{i}\prod_{k=1}^{K^{\prime}}p\left(\mathcal{R}^{k}\left(\{e|t_{i}<t(e)\leq t_{i}+\Delta T\}\right)|\mathcal{M}_{\bm{\theta}}\left(\{e_{j}\}_{j\leq i}\right),\bm{\theta}^{\prime}_{k}\right).$ (7)
We aim for the model to understand inherent motion patterns and beyond just memorize the history by learning to predict the future. In particular, individual events are ineffective prediction targets due to their limited information and unpredictable noise. In contrast, converted representations are more informative and noise-robust, offering a more reliable supervisory signal.
@@ 4.1 Object recognition
Datasets. To compare with previous A2S work martin2024alert , we evaluate our work EVA on two event-based recognition tasks: action recognition (DVS128-Gesture amir2017low ) and binary classification (N-Cars sironi2018hats ). DVS128-Gesture provides 11 classes of human actions recorded by the DVS128 event camera, containing 1342 files of around 6 seconds from 29 subjects under 3 different lighting conditions. N-Cars provides a benchmark for the binary classification task of car recognition, containing 24029 files of 100 ms recorded by the ATIS event camera.
Implementations. We present an asynchronous encoder composed of an event embedding with $D=128$, 3 RWKV-6 blocks of width $D=128$, and one MVHS layer as output to generate the learned representation. We use patch size $P=16$. On DVS128-Gesture, to reduce the latency of downstream classification, we use $D_{\text{head}}=8$ for a 2$\times$ shrink in height and width of the representation and use a light-weighted 14 layer ResNet he2016deep (referred to as ResNet-14) for classification. On N-Cars however, since the samples are already cropped small sironi2018hats , we also present an encoder with $D_{\text{head}}=16$ for a full height and weight representation. To maintain the representation size, we use only half-channels of the MVHS output. The model implementation details are given in Appendix C.
@@
On DVS128-Gesture, the encoder is pretrained on event sequences of length 8192 with batch size 32 and Adam optimizer kingma2014adam with learning rate 0.001. It takes around 10 hours to train 100 epochs with an RTX 3090 GPU. We use EC and TS for MRP and EC for NRP as SSL targets. Calculating the loss of the predictions and targets at every individual event would demand prohibitive GPU memory. To mitigate this, we divide the event sequence into chunks of length $T_{\text{chunk}}=512$, and calculate the loss only at the end of each chunk. On N-Cars, the encoder is pretrained with a sequence length of 512 and chunk length 16 with batch size 32 and learning rate 0.001. It takes 1 hour to train 20 epochs on an RTX 3090 GPU. We use EC and TS for MRP and EC for NRP as well. To balance the multi-task losses, we use uncertainty weights kendall2018multi . The details of SSL objectives are given in Appendix D.
@@
Metrics. On DVS128-Gesture, following previous work martin2024alert , we evaluate the model on samples of length 8192 and report the accuracy on samples (sample accuracy, SA) and the accuracy after majority voting on all the samples of each file (file voting accuracy, FVA). We also report the complexity (MAC) and inference latency on the RTX 3090 GPU. On N-Cars, we report accuracy.
@@
Results. On DVS128-Gesture, our model achieves 96.9% FVA and 92.9% SA (see Table 1), outperforming the best result in previous A2S work (94.1% FVA and 84.6% SA) by 2.8% and 8.3% respectively. Compared with ALERT-Tr. (+RM), our EVA achieves better performance with smaller parameter size and complexity. ALERT-Tr. (+LMM) indeed has fewer parameters, but with a much lower accuracy of 72.6% in SA, 20% less accuracy than ours. It is worth noting that our EVA, with a lightweight ResNet classifier, has much lower inference time (1.5 ms) compared with both the LMM and RM. Our asynchronous encoder has a larger inference latency compared with previous work because of its complicated architecture. However, it is smaller than the duration time of an 8192-length sample, which is about 100 ms in DVS128-Gesture. On N-Cars, previous A2S work martin2024alert did not report the results with their larger model (RM), only gave the accuracy of LMM to show its low complexity. Therefore, here we also compare our EVA with (i) synchronous tensor-based methodspeng2023get ; klenk2024masked and (ii) methods that learns the representation from raw eventsgehrig2019end ; cannici2020differentiable . In comparison with ALERT-Tr. (+LMM), our method has better accuracy and less classifier latency. Since the files in N-Cars are small in number of events and the pretraining on this dataset is not sufficient, we also present the result with representations generated by an encoder pretrained on Gen1 for detection tasks (EVA-L in Table 3). For a fair comparison with gehrig2019end and cannici2020differentiable , we follow their practice and use ResNet-34 pretrained on ImageNetdeng2009imagenet as the classifier. Our model, with the encoder pretrained on Gen1 and ResNet-34, has an accuracy of 96.3%, higher than the methods with learned representations.
@@
Table 1: Results on DVS128-Gesture. Latency refers to the time to process a sample of length 8192.
@@
Table 2: Results on N-Cars. EVA (+ResNet-14) is pretrained on N-Cars, and EVA (+ResNet-34) is pretrained on Gen1.
@@
GET peng2023get Event count $\rightarrow$ Token 96.9%
MEM klenk2024masked Event count 98.4%
EST gehrig2019end Learned 92.5%
Matrix-LSTM cannici2020differentiable Learned 95.8%
@@ 4.2 Object detection
Dataset. We evaluate our work on the challenging automotive detection dataset Gen1 de2020large , which includes 39 hours of data recorded by the Gen1 event camera with a sensor resolution of $304\times 240$.
Implementations. We present EVA of 3 RWKV-6 blocks, one MVHS layer with $D=128$. We also present a larger model with $D=192$ (referred to as EVA-L). We use patch size $P=16$ and $D_{\text{head}}=16$ to keep the spatial resolution. Like N-Cars, we use only half of the output channels of the MVHS layer. For the downstream detection task, we use RVT-B gehrig2023recurrent , a state-of-the-art (SOTA) Transformer-based synchronous backbone and YOLOX detection head yolox2021 on the learned representations. The model implementation details are given in Appendix C.
We pretrain the encoder with sequence length $T=2048$, chunk length $T_{\text{chunk}}=16$ and batch size 64 on 4 RTX 3090 GPUs. The learning rate is initially 0.001 with a decay of $\gamma=0.9$ every epoch. We use EC and TS for MRP objectives and EC for NRP. The details are given in Appendix D.
Results. Existing A2S work did not report results on object detection tasks. Therefore, here we compare our method with (i) synchronous tensor-based methods and (ii) end-to-end asynchronous methods. The results are shown in Table 3. End-to-end asynchronous methods could predict results in an event-by-event manner, but have a lower accuracy compared with synchronous methods. Our best model achieves a remarkable 47.7 mAP, and improves the SOTA synchronous method gehrig2023recurrent with a lower number of channels of the event representation (6 for our EVA compared with 20 for RVT-B). This is the first result by A2S methods on event-based detection tasks, demonstrating the effectiveness of our method.
Table 3: Object detection performance on Gen1. Async. refers to end-to-end asynchronous methods. Sync. refers to synchronous tensor-based methods.
@@
NVS-S li2021graph Async. 8.6
AEGNN schaefer2022aegnn Async. 14.5
Asynet messikommer2020event Async. 16.3
FARSE-CNN santambrogio2024farse Async. 30.0
ASTMNet li2022asynchronous Sync. 46.7
HMNet-L3 hamaguchi2023hierarchical Sync. 47.1
RVT-B gehrig2023recurrent Sync. 47.2
GET peng2023get Sync. 47.9
@@ 4.3 Ablation study
We ablate architecture components, including MVHS and temporal embedding, and SSL objectives to check their contribution. The effect of patch size is also investigated. The ablation is conducted on DVS128-Gesture, following the implementations in the object recognition experiment. We report the loss items of SSL during pretraining and downstream task accuracy, including FVA and SA, on the validation set.
@@
0.366 3.16×$\bf{\times 10^{-3}}$ 0.767 98.1% 94.7%
1.870 4.56×$\times 10^{-2}$ 0.933 87.8% 81.1%
0.826 9.34×$\times 10^{-3}$ 0.834 97.4% 94.1%
Ablation of SSL objective. We remove the SSL objectives to show their contribution to the learned representation. Results in Table 5 show that removing the items of SSL decreases the performance of the model. One interesting thing is that learning only one representation, such as EC, will not contribute to a smaller pretraining loss compared with learning all the representations. This indicates that learning one single representation could benefit from learning other representations.
@@ 5 Conclusion
We introduce a new A2S framework, EVA, for real-time event camera data processing. The framework consists of an efficient and expressive linear attention-based asynchronous encoder for event-by-event representation updating and an SSL method to learn a generalizable event representation. EVA outperforms the previous A2S work martin2024alert , achieves remarkable performance on event-based detection tasks, and learns generalizable representations applicable in diverse downstream tasks, demonstrating its potential towards real-time event-based applications.
The limitations of our work are discussed in Appendix G.
Appendix A Architecture details of the asynchronous encoder
RWKV-6 details. As described in [33], one RWKV-6 block is composed of token mixing (TM) and channel mixing (CM). For input sequence $\{\bm{x}_{i}\in\mathbb{R}^{D}\}_{i=1}^{T}$ of length $T$, TM first conduct a data-dependent linear interpolation (ddlerp) between $\bm{x}_{i}$ and $\bm{x}_{i-1}$ by

### v2

@@ 3.2.1 Multi-representation prediction
For MRP, we force the encoded feature to predict multiple handcrafted (converted) event representations (Lagorce et al., 2016; Sironi et al., 2018; Innocenti et al., 2021), such as event count (EC) and time surface (TS), to learn a more comprehensive feature. For events $\mathcal{E}=\{e_{i}=(t_{i},x_{i},y_{i},p_{i}),i=1,2,\dots\}$, we expect the asynchronous encoder $\mathcal{M}_{\bm{\theta}}$ parametrized by $\bm{\theta}$ to learn an event-by-event feature as $\mathcal{F}_{i}=\mathcal{M}_{\bm{\theta}}(\{e_{j}\}_{j\leq i})$. Given $K$ handcrafted representations $\{\mathcal{R}^{k}\}_{k=1}^{K}$, we predict these representations from $\mathcal{F}$ with $K$ specific heads parameterized by $\bm{\Theta}=\{\bm{\theta}_{k}\}_{k=1}^{K}$. The objective is
$\arg\max_{\bm{\theta},\bm{\Theta}}\mathbb{E}_{i}\prod_{k=1}^{K}\mathbf{Pr}\left(\mathcal{R}^{k}_{i}|\mathcal{F}_{i}=\mathcal{M}_{\bm{\theta}}\left(\{e_{j}\}_{j\leq i}\right);\bm{\theta}_{k}\right),$ (4)
where $\mathbf{Pr}$ means probability. MRP is motivated by the insight that different handcrafted representations capture different aspects of information about the raw events. They provide the required information for various tasks. Therefore, we anticipate that learning from multiple converted representations will yield a feature generalizable to diverse downstream tasks.
@@ 3.2.2 Next representation prediction
Inspired by auto-regressive pretraining of language models, where models learn by next-token prediction (Radford et al., 2018), NRP forces the encoded feature to predict handcrafted representations for a future time interval. For a future time window $\Delta T$, the objective of NRP is
$\arg\max_{\bm{\theta},\bm{\Theta}^{\prime}}\mathbb{E}_{i}\prod_{k=1}^{K^{\prime}}\mathbf{Pr}\left(\mathcal{R}^{k}\left(\{e|t_{i}<t(e)\leq t_{i}+\Delta T\}\right)|\mathcal{F}_{i}=\mathcal{M}_{\bm{\theta}}\left(\{e_{j}\}_{j\leq i}\right);\bm{\theta}^{\prime}_{k}\right).$ (5)
We aim for the model to understand inherent motion patterns and go beyond just memorizing the history by learning to predict the future. In particular, individual events are ineffective as prediction targets due to their limited information and unpredictable noise. In contrast, converted representations are more informative and noise-robust, offering a more reliable supervision signal.
@@ 4.1 Object recognition
Datasets. To compare with previous A2S work (Martin-Turrero et al., 2024), we evaluate our EVA on two event-based recognition tasks: action recognition (DVS128-Gesture (Amir et al., 2017)) and binary classification (N-Cars (Sironi et al., 2018)). DVS128-Gesture provides human actions recorded by the DVS128 event camera, containing 1342 files of around 6 seconds from 29 subjects under 3 different lighting conditions. N-Cars provides a benchmark for the binary classification task of car recognition, containing 24029 files of 100 ms recorded by the ATIS event camera.
[…]
@@
On DVS128-Gesture, the encoder is pretrained on event sequences of length 8192 with batch size 32 and Adam optimizer (Kingma and Ba, 2014) with learning rate 0.001. It takes 10 hours to train 100 epochs with an RTX 3090 GPU. On N-Cars, the encoder is pretrained with a sequence length of 512 with batch size 32 and learning rate 0.001. It takes 1 hour to train 20 epochs on an RTX 3090 GPU. We use EC and TS for MRP and EC for NRP as the target. To balance the multi-task losses, we use uncertainty weights (Kendall et al., 2018). The details of loss items are given in Appendix E.
@@
Metrics. On DVS128-Gesture, following previous work (Martin-Turrero et al., 2024), we evaluate the model on samples of length 8192 and report the accuracy on samples (sample accuracy, SA) and the accuracy after majority voting on all the samples of each file (file voting accuracy, FVA). We also report the complexity (MAC) and inference latency on the RTX 3090 GPU. On N-Cars, we report accuracy.
@@
Results. On DVS128-Gesture, our model achieves 96.9% FVA and 92.9% SA (see Table 1), outperforming the best result in previous A2S work (94.1% FVA and 84.6% SA) by 2.8% and 8.3% respectively. Compared with ALERT-Tr. (+RM), our EVA achieves better performance with a smaller parameter size and complexity. ALERT-Tr. (+LMM) indeed has fewer parameters, but with a much lower accuracy of 72.6% in SA, 20% less accuracy than ours. It is worth noting that our EVA, with a lightweight ResNet classifier, has much lower inference time (1.5 ms) compared with both the LMM and RM. Our asynchronous encoder has a larger inference latency compared with previous work because of its hierarchy learning architecture. However, it has a shorter processing time for an 8192-length sample, which is about 100 ms in DVS128-Gesture. On N-Cars (Table 2), previous A2S work (Martin-Turrero et al., 2024) did not report the results with their best model (RM), only gave the accuracy of LMM to show its low complexity. Therefore, here we additionally compare our EVA with (i) dense event image-based methods (Peng et al., 2023b; Klenk et al., 2024) and (ii) methods that learn the representation/feature from raw events (Gehrig et al., 2019; Cannici et al., 2020). In comparison with ALERT-Tr. (+LMM), our method has better accuracy and less classifier latency. Since the files in N-Cars are small in number of events and the pretraining on this dataset is not sufficient, we also present the result with features generated by an encoder pretrained on Gen1 for detection tasks (EVA-L in Table 3). For a fair comparison with Gehrig et al. (2019) and Cannici et al. (2020), we follow their practice and use ResNet-34 pretrained on ImageNet (Deng et al., 2009) as the classifier. Our model, with the encoder pretrained on Gen1 and ResNet-34, has an accuracy of 96.3%, higher than the methods with learned representations.
@@
Table 1: Results on DVS128-Gesture. Latency refers to the time to process a sample of length 8192. We compare with previous A2S work ALERT-Tr. (Martin-Turrero et al., 2024).
@@
Table 2: Results on N-Cars. EVA is pretrained on N-Cars, and EVA-L is pretrained on Gen1.
@@
GET (Peng et al., 2023b) Event count $\rightarrow$ Token 96.9%
MEM (Klenk et al., 2024) Event count 98.4%
EST (Gehrig et al., 2019) Learned 92.5%
Matrix-LSTM (Cannici et al., 2020) Learned 95.8%
@@ 4.2 Object detection
Dataset. We evaluate our work on the challenging automotive detection dataset Gen1 (De Tournemire et al., 2020), which includes 39 hours of data recorded by the Gen1 event camera with a sensor resolution of $304\times 240$.
Implementations. We present EVA of 3 RWKV-6 blocks, one MVHS layer with $D=128$. We also present a larger model with $D=192$ (referred to as EVA-L). We use patch size $P=16$ and $D_{\text{head}}=16$ to keep the spatial resolution. Like N-Cars, we use only half of the output channels of the MVHS layer. For the downstream detection task, we use RVT-B (Gehrig and Scaramuzza, 2023), a lightweight state-of-the-art (SOTA) Transformer-based synchronous backbone and YOLOX detection head (Ge et al., 2021) on the learned features. The model implementation details are given in Appendix D.
We pretrain the encoder with sequence length $T=2048$ and batch size 64 on 4 RTX 3090 GPUs. The learning rate is initially 0.001 with a decay of $\gamma=0.9$ every epoch. We use EC and TS for MRP objectives and EC for NRP. The details are given in Appendix E.
Results. Existing A2S work did not report results on object detection tasks. Therefore, here we compare our method with (i) synchronous dense methods and (ii) end-to-end asynchronous methods. The results are shown in Table 3. Asynchronous methods have minimal latency but lower accuracy compared with synchronous methods. Our best model achieves a 47.7 mAP, and improves the SOTA synchronous method (Gehrig and Scaramuzza, 2023) with a lower number of channels of the input feature (6 for our EVA compared with 20 for RVT-B). This is the first result by A2S methods on event-based detection tasks, demonstrating the effectiveness of our method.
Table 3: Object detection performance on Gen1. A refers to end-to-end asynchronous methods. S refers to synchronous dense methods.
@@
NVS-S (Li et al., 2021) A 8.6
AEGNN (Schaefer et al., 2022) A 14.5
Asynet (Messikommer et al., 2020) A 16.3
FARSE-CNN (Santambrogio et al., 2024) A 30.0
DAGr-L (Gehrig and Scaramuzza, 2024) A 32.1
ASTMNet (Li et al., 2022) S 46.7
HMNet-L3 (Hamaguchi et al., 2023) S 47.1
RVT-B (Gehrig and Scaramuzza, 2023) S 47.2
GET (Peng et al., 2023b) S 47.9
@@
4.3 Timing experiments
MACs. per event. We timed our EVA models, implemented in Python and CUDA, on an RTX 3090 GPU. We summarize the results in Table 4 and the event rate (number of events per second) of the datasets in Table 5. For DVS128-Gesture and N-Cars, the event rates are both within 100,000 per second, lower than the throughput of the EVA models (at least 541,000 per second). For Gen1 with a larger resolution, the event rate surpasses the throughput of our EVA-L model. Since patch-wise encoding (PWE) is used in our EVA encoder, the features of each patch could be encoded independently. Therefore, considering that the patch event rate of the Gen1 dataset is 2.17 K/s, much lower than the throughput of EVA-L, the real-time encoding on Gen1 could still be achieved.
Table 4: Throughput and MACs of EVA models.
Model # Param MACs per Event Throughput (M/s)
EVA 0.6 M 0.60 M 0.611
EVA($D_{head}=16$) 0.6 M 0.61 M 0.560
EVA-L 1.4 M 1.31 M 0.541
Table 5: Event rates of different datasets.
Dataset Camera Resolution Event rate (M/s) Patch event rate (K/s)
DVS128-Gesture DVS128 $\times 128$ 0.058 0.90
N-Cars ATIS $0×\times 1$20 0.039 4.07
[…]
Latency vs. Context length. We evaluate the inference time with respect to input context length. As shown in Figure 5, the inference time is basically proportional to the context length (number of input events). We could leverage temporal subsampling (Santambrogio et al., 2024) or pooling (Soydan et al., 2024) to reduce MACs, which are commonly used in asynchronous event processing.
Figure 5: Inference latency on different sequence lengths.
Latency vs. Sensor resolution. For high-resolution event cameras (i.e., Gen 3 with 1280 $\times$ 720 pixels), the event rate would increase proportionally to the sensor size. Therefore, these high-frequency event sequences would challenge our EVA encoder. Theoretically, since EVA conducts a PWE strategy, the features of each patch could be computed in parallel. The overall runtime is only determined by the longest computation time among all the patches, and thus only determined by the patch event rate. From Table 5, the patch event rate is stable across event cameras with different resolutions. Therefore, our method could, in principle, be applicable to event cameras at any resolution, without introducing further latency.
@@ 4.4 Ablation study
We ablate architecture components (MVHS and temporal embedding) and SSL objectives to check their contribution. The effect of patch size is also investigated. The ablation is conducted on DVS128-Gesture, following the implementations in the object recognition experiment. We report the loss items of SSL during pretraining and downstream accuracy on the validation set.
@@
[…]
Ablation of SSL objective. We remove the SSL objectives to show their contribution to the learned feature. Results in Table 7 show that removing the items of SSL decreases the performance of the model. One interesting thing is that learning only one representation, such as EC, will not contribute to a smaller pretraining loss compared with learning all the representations. This indicates that learning one single representation could benefit from learning other representations.
@@ 5 Conclusion
We introduce a new A2S framework, EVA, for real-time event camera data processing. The framework consists of an efficient and expressive linear attention-based asynchronous encoder for event-by-event feature updating and an SSL method to learn a generalizable event feature. EVA outperforms the previous A2S work (Martin-Turrero et al., 2024), achieves remarkable performance on event-based detection tasks, and learns generalizable features applicable in diverse downstream tasks, demonstrating its potential towards real-time event-based applications.
The limitations of our work are discussed in Appendix H.
Reproducibility statement
Code: Code is available https://github.com/haohq19/eva. Datasets: The data preprocessing steps are described in Appendix C. Experiments: All experimental settings, including hyperparameters and model architectures, are detailed in Section 4 and Appendix D. The random seeds used for all experiments are fixed.
Acknowledgement
This work was supported by STI 2030-Major Projects 2021ZD0200300 and High Performance Computing Center of Tsinghua University.
Appendix A Extended experimental results
A.1 Object recognition
We provide object recognition experiments on static-captured event camera datasets. We evaluate the EVA feature on N-Caltech101 (Orchard et al., 2015) dataset, which is created by moving an event camera that is focused on a screen displaying the original Caltech101 dataset samples. The dataset consists of 100 object classes and an additional background class. Since the dataset does not provide an official train/test split, we follow the split in EST (Gehrig et al., 2019) for fair comparison. The results are shown in Table 9. Our EVA method outperforms all the asynchronous methods and has competitive accuracy compared with synchronous methods. Our method surpasses EST (Gehrig et al., 2019) and MatrixLSTM (Cannici et al., 2020), which learn representations in a supervised, end-to-end manner. E2VID reconstructs gray-scale images and introduces an extra latency (Rebecq et al., 2019). Our EVA method, however, does not lead to further latency and only shows a marginal 0.3% drop in accuracy. This demonstrates that our EVA feature could also be used on the challenging static-captured datasets.
Table 9: Object recognition results on N-Caltech101. A. refers to asynchronous. S. refers to synchronous.
Model Type Acc.
HOTS (Lagorce et al., 2016) A. 21.0 %
NVS-S (Li et al., 2021) A. 67.0 %
AEGNN (Schaefer et al., 2022) A. 66.8 %
Asynet (Cannici et al., 2019) A. 74.5 %
FARSE-CNN (Santambrogio et al., 2024) A. 68.7 %
EST (Gehrig et al., 2019) S. 83.7 %
Matrix-LSTM (Cannici et al., 2020) S. 84.3 %
Pillar (Fan et al., 2025) S. 85.3 %
E2VID (Rebecq et al., 2019) S. 86.6 %
EVA-L (ours) A2S 86.3 %
A.2 Object detection
To better demonstrate the expressivity and low-latency advantage of the learned asynchronous EVA feature, here we compare our EVA method with other event-representing methods under an identical neural network backbone (Swin Transformer V2). We compare both accuracy and representing latency, which refers to the time required for computing the representation when a new event arrives (Fan et al., 2025). The results are shown in Table 10.
Table 10: Comparison of different event representations with an identical Swin V2 backbone on the Gen1 detection task. Representing latency refers to the time required for computing the representation when a new event arrives.
Method mAP (%) Asynchronous Representing latency (ms)
Voxel(Fan et al., 2025) 39.5 16.3
ERGO-12(Zubić et al., 2023) 50.4 >> 10
Pillar(Fan et al., 2025) 53.1 11.4
HOTS(Lagorce et al., 2016) 49.0 1.0
TORE(Baldwin et al., 2022) 43.6 << 0.1
EVA-L (ours) 49.7 << 0.1
EVA (ours) 49.0 << 0.1
Synchronous representations provide better performance, but with additional latency. Our EVA method outperforms all the asynchronous representations and maintains its low-latency characteristic. This further demonstrates the advantages of our method in low-latency applications.
A.3 Timing experiments
[…]
Table 11: Accuracy and throughput under different event rates with downsampling.
Model Length SA. FVA. Throughput (K/s)
EVA (ours) 8192 92.9 96.9 611
4096 91.8 96.2 1222
2048 91.3 95.3 2444
1024 90.0 95.1 4888
512 88.4 93.4 9776
ALERT-Tr. (+RM) 8192 84.6 94.1 1412
ALERT-Tr. (+LMM) 8192 72.6 86.2 2100
Appendix B Architecture details of the asynchronous encoder
RWKV-6 details. As described in (Peng et al., 2024), one RWKV-6 block is composed of token mixing (TM) and channel mixing (CM). For input sequence $\{\bm{x}_{i}\in\mathbb{R}^{D}\}_{i=1}^{T}$ of length $T$, TM first conduct a data-dependent linear interpolation (ddlerp) between $\bm{x}_{i}$ and $\bm{x}_{i-1}$ by